# Enhancing Lip Reading with Multi-Scale Video and Multi-Encoder

### v1

@@ 3.4 Ablation Study
To prove the effectiveness of our proposed Enhanced ResNet3D visual front-end, data augmentation, speed perturbation, and fine-tuning on the performance of ALR systems, we conduct ablation experiments based on the Branchformer encoder using 1.5 scale video data. The ablation experiment results are shown in Table 2. A0 is the ALR system based on the ResNet3D visual front-end and Branchformer proposed by Wang et al. [17], which we reproduce on the ChatCLR challenge dataset. While A1 replaces the visual front-end with our proposed Enhanced ResNet3D, yielding a 1.48% reduction in CER on the evaluation set. It is noteworthy that, comparing the results of systems A2 and A1, we find that incorporating the random data augmentation strategy during training yields significant benefits, with a 3.61% reduction in CER on the evaluation set. Additionally, with speed perturbation on the training set, system A3 achieves a 1.4% CER reduction compared to A2. Finally, after fine-tuned with training set data from the same speakers as the evaluation set and 90% of the development set data, system S11 achieves the best single-system performance, further reducing the CER on the evaluation set by 0.88% compared to system A3.
@@ 4 Conclusion
This paper summarizes our work in the target speaker lipreading task of the ICME2024 Chat-scenario Chinese Lip Reading (ChatCLR) challenge. We propose a novel lip-reading approach that uses multi-scale lip motion video data and different visual encoders to build diverse automatic lip-reading (ALR) systems, and performs multi-system fusion on all output transcripts. For the ALR system, we propose an Enhanced ResNet3D visual front-end (VFE) module for extracting visual features, and introduce Branchformer and E-Branchformer as the visual encoder. We report the results of all ALR systems and analyze three aspects: data, visual encoders, and multi-system fusion. Ablation experiments are also conducted to show the effectiveness of the Enhanced ResNet3D VFE, data augmentation, speed perturbation, and fine-tuning. Finally, we achieve a 78.17% character error rate (CER) on the challenge evaluation set with a 21.52% reduction compared to the official baseline, ranking second place.

### v2

@@ 3.4 Ablation Study
To prove the effectiveness of our proposed Enhanced ResNet3D visual front-end, data augmentation, speed perturbation, and fine-tuning on the performance of ALR systems, we conduct ablation experiments based on the Branchformer encoder using 1.5 scale video data. The ablation experiment results are shown in Table 2. A0 is the ALR system based on the ResNet3D visual front-end and Branchformer proposed by Wang et al. [17], which we reproduce on the ChatCLR challenge dataset. While A1 replaces the visual front-end with our proposed Enhanced ResNet3D, yielding a 1.48% reduction in CER on the evaluation set. It is noteworthy that, comparing the results of systems A2 and A1, we find that incorporating the random data augmentation strategy during training yields significant benefits, with a 3.61% reduction in CER on the evaluation set. Additionally, with speed perturbation on the training set, system A3 achieves a 1.4% CER reduction compared to A2. Finally, after fine-tuning with training set data from the same speakers as the evaluation set and 90% of the development set data, system S14 achieves the best single-system performance, further reducing the CER on the evaluation set by 0.88% compared to system A3.
@@ 4 Conclusion
This paper summarizes our work in the target speaker lipreading task of the ICME2024 Chat-scenario Chinese Lip Reading (ChatCLR) challenge. We propose a novel lip-reading approach that uses multi-scale lip motion video data and different visual encoders to build diverse automatic lip-reading (ALR) systems, and performs multi-system fusion on all output transcripts. For the ALR system, we propose an Enhanced ResNet3D visual front-end (VFE) module for extracting visual features, and introduce Branchformer and E-Branchformer as the visual encoder. We report the results of all ALR systems and analyze three aspects: data, visual encoders, and multi-system fusion. Ablation experiments are also conducted to show the effectiveness of the Enhanced ResNet3D VFE, data augmentation, speed perturbation, and fine-tuning. Finally, we achieve a 78.17% character error rate (CER) on the evaluation set with a 21.52% reduction compared to the official baseline, ranking second place.